OpenStudy (hang254):

Find an equation for the nth term of the arithmetic sequence. -13, -8, -3, 2, ...

OpenStudy (hang254):

@jim_thompson5910 @campbell_st @paki

Mehek (mehek14):

how do you get -8 from -13?

OpenStudy (hang254):

add 5

OpenStudy (hang254):

@Mehek14

Mehek (mehek14):

ok from -8 to -3?

OpenStudy (hang254):

add 5 again

Mehek (mehek14):

so 5 is your constant rate of change now can you make an equation?

OpenStudy (hang254):

\[a _{n}=-13+5(n+1)\]

OpenStudy (hang254):

is that correct?

Mehek (mehek14):

should it be \(\large{n - 1}\)?

OpenStudy (hang254):

how come n-1 ?

Mehek (mehek14):

isn't the origional one \(\large{a_n+1 = a_n + 5}\)

Mehek (mehek14):

then you move it to the other side it becomes -1

OpenStudy (hang254):

ohh thats right, i see it now

OpenStudy (hang254):

so \[a _{n}=-13 +5(n-1)\]

Mehek (mehek14):

so it would be \(\Large{a_n = -11 + 5(n-1)}\) Then you use distributive property

OpenStudy (hang254):

thank you

Mehek (mehek14):

your welcome